A 25.00 mL sample of 0.379 M nitric acid (HNO3(aq)) is titrated with 12.00 mL of 0.573 M lithium hydroxide (LiOH(aq)). Determine the amount of [H+(aq)] remaining in the flask and pH of the solution after reaction.
1
Expert's answer
2019-08-22T03:21:34-0400

Solution.

"LiOH + HNO3 = LiNO3 + H2O"

"OH^{-} + H^{+} = H2O"

"C = \frac{n}{V}"

"n(LiOH) = C(LiOH) \times V(LiOH)"

"n(HNO3) = C(HNO3) \times V(HNO3)"

n(LiOH) = 0.006876 mole

n(HNO3) = 0.009475 mole

n(LiOH)<n(HNO3), therefore lithium hydroxide will be limiting reagent.

Unconsumed nitric acid remained in the solution, so pH will be less than 7.

"n(HNO3 (left)) = 0.009475-0.006876 = 0.002599 mole"

"[H^{+}(aq)] = n(HNO3 (left)) = 0.002599 mole"

"pH = -lg([n(H^{+})])"

pH = 2.59

Answer:

"[H^{+} (aq)] = 0.002599 mole"

"pH = 2.59"